OpenStudy (across):

Fun DE time! Given\[f(x)=\begin{cases} -1 & \text{ if } 0\leq t<1 \\ 1 & \text{ if } 0\geq1 \end{cases}\]Find\[\mathcal{L}\{f(x)\}\]

OpenStudy (pdoctor):

Drats, I don't remember.

OpenStudy (across):

\[\mathcal{L}\{f(t)\}=\int_{0}^{1}e^{-st}(-1)dt+\int_{1}^{\infty}e^{-st}(1)dt.\]

OpenStudy (jamesj):

Let u(t) = 1, x >= 0 0, x < 0 the unit step function. Then your function f(x) = -2u(t) + 3u(t-1) and the Laplace transform is just the linear combination of these L{f(x)} = -2 L{u(t)} + 3 L{u(t-1)} = -2/s + 3e^(-s)/s

OpenStudy (across):

The unit step function... ... I somehow got\[\mathcal{L}\{f(t)\}=\frac{2}{s}e^{-s}-\frac{1}{s}.\]I'll recheck.

OpenStudy (pdoctor):

You got the right answer.

OpenStudy (jamesj):

oh, sorry, I wrote down f incorrectly f = -u(t) + 2u(t-1) So L(f(t)) = -1/s + 2e^(-s)/s the same thing

OpenStudy (jamesj):

Hate it when if forget things like \[-2 \neq -1\]